Circular No. 5932

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

γ-RAY BURST A. Goldwurm, S. Gaertner, E. Churazov, and M. Gilfanov, on behalf

of the Granat SIGMA team (Service d’Astrophysique, Centre d’Etudes de Saclay; Centre d’Etude Spatiale des Rayonnements, Toulouse; Space Research Institute, Moscow), report: “The SIGMA telescope onboard the Granat Observatory detected a γ-ray burst on Feb. 6.00625 UT, which lasted ∼ 16 s and was detected in the energy range 40-480 keV through an interruption of the passive shield acting as a collimator. The localization process is based on a maximum likelihood method, yielding a signal-to-noise ratio of 24. The 1σ error box of the burst is approximately rectangularshaped, the corners having the following positions: α = 10^{h}04^{m}00^{s}, δ = −68^{o}40′.2; 9^{h}57^{m}19^{s}, −68^{o}00′.0; 10^{h}46^{m}41^{s}, −57^{o}19′.8; 11^{h}01^{m}19^{s}, −56^{o}19′.8 (equinox 1950.0).”

η CARINAE

S. White, University of Maryland; R. Duncan, Australia Telescope; S. Drake, Goddard Space Flight Center; J. Lim, California Institute of Technology; and M. Kundu, University of Maryland, report: “Australia Telescope observations of the luminous blue variable η Car show that it is undergoing a radio outburst, with the flux having doubled in the last 7 months and apparently still increasing. It is currently a 2-Jy radio source at 9 GHz. The rise in flux is associated with the appearance of a new feature in the radio images, ∼ 1′′.1 northwest of the radio source attributed to the stellar wind and almost as bright. Assuming that this feature is due to a gas clump ejected asymmetrically by the star and radiating thermal free-free emission, the minimum mass involved is 0.005 M_{⊙}. Infrared observations to follow the history of dust formation in this gas clump would be valuable.”

PERIODIC COMET SCHWASSMANN-WACHMANN 1 Total visual magnitude estimates (cf. IAUC 5915): Jan. 10.31 UT,

11.9 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 23.19, 12.3 (H. Mikuž, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + CCD + V filter); 29.7, 13.3 (Mikuž); Feb. 7.56, 13.3 (T. Kojima, Chiyoda, Ohra, Gunma, Japan, 0.25-m reflector + CCD); 10.57, 12.3 (Kojima).

1994 February 10 Daniel W. E. Green